oregonduck (oregonduck):

A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles? Enter your answer in the box. inches Part B How many miles are represented by 1 4 inch on the map? Enter your answer in the box. miles

oregonduck (oregonduck):

@Nnesha

OpenStudy (jack1):

2 inches : 25 miles = x inches : 60 miles so fractions, right?

oregonduck (oregonduck):

correct

oregonduck (oregonduck):

is it x=5?

OpenStudy (jack1):

cool, so i'll start and you solve, k? \(\huge \frac 2{25} = \frac x{60}\) \(\huge \frac {2 \times 60}{25} = \frac {x\times 60}{60}\) \(\huge \frac {120}{25} = \frac x{1}\) \(\huge \frac {120}{25} = x\)

oregonduck (oregonduck):

x= 4.8

OpenStudy (jack1):

perfect!

oregonduck (oregonduck):

so will x be the same for part b?

OpenStudy (jack1):

no, x is how many inches for 60 miles only

OpenStudy (jack1):

is that 1/4 or is it 14? for part B?

oregonduck (oregonduck):

oh...ok then it is 1/4

OpenStudy (jack1):

ok, so 1/4 = 0.25, yeah? \(\huge \frac 2{25} = \frac {0.25}{y}\) can you solve for y from here?

OpenStudy (jack1):

so if 2 inches : 25 miles , 0.25 inches : y miles, right?

oregonduck (oregonduck):

yup but how to get y still a little confused

oregonduck (oregonduck):

wait, i got it y=3.125

OpenStudy (jack1):

awesome, yep perfect dude!

oregonduck (oregonduck):

alright thxs @Jack1 for your help

OpenStudy (jack1):

np bro, slaters ;)

OpenStudy (nysha):

did you get this right?